OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 1) and (0, -3)? 1 - 1 2 -2 1 5

OpenStudy (conqueror):

Check your inbox

OpenStudy (imstuck):

Slope formula is\[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]Those x1 and x2 and y1 and y2 come from the points you provided us.

OpenStudy (igreen):

You can use the slope formula. \(m = \dfrac{y2-y1}{x2-x1}\) To find the slope between two points. Our points are: y2: -3 y1: 1 x2: 0 x1: -2

OpenStudy (anonymous):

is the 2 and 1 the exponents?

OpenStudy (igreen):

No, there aren't any exponents, I gave you the numbers, just put them in. No exponents, just regular numbers.

OpenStudy (anonymous):

-3.5

OpenStudy (igreen):

How did you get that?

OpenStudy (igreen):

\(m = \dfrac{-3-1}{0+2}\)

OpenStudy (anonymous):

ty

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @math-geek